Differentiate the following with regard to x :

y=x^4 times sin x


1
Expert's answer
2021-07-08T09:36:14-0400

Given "y=x^4 sin x"


Using chain rule, "\frac{}{}" if y= uv

"\frac{dy}{dx} = u\frac{dv}{dx}+v\frac{du}{dx}"


Then,

"\frac{dy}{dx} = x^4\frac{d sinx}{dx}+sinx\frac{dx^4}{dx}"


"\frac{dy}{dx} = x^4 cosx + 4x^3sinx"


"\frac{dy}{dx} = x^3(xcosx+4sinx)"